OpenStudy (toxicsugar22):

how do u solve log(5+x)-log(x-4)=log 2

OpenStudy (anonymous):

log a - log b = log(a/b)

OpenStudy (anonymous):

log (5+x)/(x-4) = log 2 (5 + x)/(x+4) = 2 - solve this in normal way

OpenStudy (anonymous):

Subtraction of logs becomes a division so: log(5+x)-log(x-4)=log((5+x)/(x-4))=log(2). Raise everything to the 10^ giving. (5+x)/(x-4)=2 Then multiply, combine terms, etc.

OpenStudy (anonymous):

use property of log: log b= log(a)<=>a=b loga-logb=log(a/b)

OpenStudy (toxicsugar22):

so it would be (5+x)/(x-4)=2

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

well u got to solve that for x of course

OpenStudy (toxicsugar22):

of i got 13

OpenStudy (anonymous):

Thats what I got with mental math xP

OpenStudy (toxicsugar22):

thank you

OpenStudy (anonymous):

(5+x) = 2(x-4) x = 5+8 = 13 - yes

OpenStudy (toxicsugar22):

can u help me with this one How long will it take for a poulation of 4000 to double if its annual growth is 7.1% (Round to the nearest year)

OpenStudy (anonymous):

ok 8000 = 4000(1 + (7.1/100))^x (1 + (7.1/100))^x = 2 taking logs: x log 1.071 = log 2 x = log 2/ log 1.071 x = 10.1 = 10 years

OpenStudy (toxicsugar22):

thanks i i am goint to take my test right now my final

OpenStudy (toxicsugar22):

so bye